OpenStudy (anonymous):

determine whether the expression is a polynomial. if it is, state how many terms and variables the polynomial contains. then state its degree.

OpenStudy (anonymous):

\[-5x+\frac{ 1 }{3x }\]

OpenStudy (anonymous):

Not a polynomial. For any polynomial, the degree of the variable must be an integer value greater than or equal to 0.

OpenStudy (anonymous):

ok what about this one\[-3x ^{-4}y ^{4}\]

OpenStudy (anonymous):

Here we notice that:\[\bf -5x+\frac{1}{3x}=-5x+\frac{1}{3}x^{-1}\]Clearly, the second 'x' has a degree less than one.

OpenStudy (anonymous):

Like I said, the degree of the variables must be an integer value greater than or equal to 0. Is this true for the second one? (Hint: Look at the exponent of 'x')

OpenStudy (anonymous):

no it is less than 0

OpenStudy (anonymous):

so this one would not be a polynomial either

OpenStudy (anonymous):

exactly.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

For a better understanding of Polynomials, I really think this page will help you out in a fun and simple way: http://www.mathsisfun.com/algebra/polynomials.html @cahayes9498